A sidewalk vendor has 3 items on sale namely, a pocketbook for $50,a wall clock at $30, and a pencil at $1. At the end of the day he has sold a total of 100 items for a total sale of $1000. How many pocketbooks were sold?
Answer by JBarnum(2146)

50b+30w+p=1000
b+w+p=100

not enough info so imma guess n check
with looking at top equation if the were no wall clocks sold or pencils the most amount of pocketbooks couls be 1000/50=20, same procedure for wall clocks is 1000/30=33, pencils would be 1000.
so we know that theres: b<20 w<30 and p<1000
looking at second equation p<100
b w p
10+10+80=100
500+300+80=880
wrong
5+5+90=100
250+150+90=490
wrong
20+30+50=100
1000+900+50=1950
wrong
15+5+80=100
750+150+80=980
wrong
16+4+80=100
800+120+80=1000
correct.....yes found it
16 pocketbooks 4 wall clocks and 80 pencils sold